If f(x) = x-3/5, then the value of f(8) equals...: If f(x) = x-3/5, then the value of f(8) equals...
Answer by oscargut(667)

If f(x) = x-3/5, then the value of f(8)is:
f(8)=8-3/5=40/5-3/5=37/5
Answer: 37/5